OpenStudy (anonymous):

Multiply the equation by a power of 10 to write an equivalent equation with integer coefficients. 4.603y - 1.842 = -3.651y

OpenStudy (mathstudent55):

Notice that here all decimals have three decimal places. That means the thousandths place.That means you need to multiply every one of the numbers by 1000.

OpenStudy (anonymous):

I don't understand

OpenStudy (mathstudent55):

When you multiply a number by ten, you move the decimal place one place tot he right. 0.5 x 10 = 5 0.8 x 10 = 9 0.23 x 10 = 2.3

OpenStudy (mathstudent55):

If you multiply a number by 100, you move the decimal point two places to the right. 0.05 x 100 = 5 0.067 x 100 = 6.7 5.678 x 100 = 567.8

OpenStudy (mathstudent55):

When you multiply a number by 1000, you move the decimal point 3 places to the right. 0.006 x 1000 = 6 25.678 x 1000 = 25,678

OpenStudy (mathstudent55):

In this problem you have three numbers that have three places after the decimal point. The numbers are 4.603, 1.842, and 3.651. Your goal is to make whole numbers out of them. Since they have 3 places after the decimal point, you need to multiply them by 1000 because that will make decimal point move three places to the right, and they will no longer have a decimal part. 4.603 x 1000 = 4603 1.842 x 1000 = 1842 3.651 x 1000 = 3651

OpenStudy (mathstudent55):

Powers of 10 are 10^1 = 10 10^2 = 100 10^3 = 1000 10^4 = 10,000 etc. and also 10^0 = 1 10^(-1) = 1/10 = 0.1 10^(-2) = 1/100 = 0.01 10^(-3) = 1/1000 = 0.001 etc.

OpenStudy (anonymous):

Okay so a power of 10 is like for instance 2^10, 2^100, 2^1000 etc. The powers of 10 go on with the more zeroes you add. In this problem, they want you to make the numbers with decimals into integers, so that means we have to make them lose their decimal points. By every 10th power we multiply one of those decimal numbers the decimal point will move to the right. .562*10=5.62 .562*100=56.2 .562*1000=562 Now if we multiply each number in the problem above by 1000, they'll become integers. First we'll start with 4.603 4.603*10=46.03 4.603*100=460.3 4.603*1000=4603 With every zero we add to the power, the decimal place moves one more spot to the right. Now we'll multiply those powers with 1.842- 1.842*10=18.42 1.842*100=184.2 1.842*1000=1842 Finally we'll multiply those powers with 3.651 3.651*10= 36.51 3.651*100=365.1 3.651*1000=3651 Now we pick the numbers that we multiplied by 1000 because the became integers as the decimal place left. So our answer is- 4603y-1842=-3651y

OpenStudy (anonymous):

thanks ya'll ^-^

OpenStudy (anonymous):

no prob. :)